Put the function y=10x(x+6) in factored form f(x)=a(x-r)(x-s) and state the values of a, r, and s.
a=
r=
s=

Answer by Theo(13342)

i'm not exactly sure what they are looking for.

the equation looks like it is already in factored form.

it appears to be 10 * (x+6) * (x + 0).

the form they are looking for is a * (x-r) * (x-s)

if i'm guessing correctly, then you would get:

a = 10
r = -6
s = 0

the unfactored equation would be 10x^2 + 60x = 0

you would factor out the gcf to get 10 * (x^2 + 6x)

if you tried to factor the quadrtic of x^2 + 6x using the quadratic formula, you would get:

a = 1
b = 6
c = 0

sqrt(b^2 - 4ac) = sqrt(36-4*1*0) = sqrt(36) = 6

quadratic formula of x = (-b +/- sqrt(b^2-4ac))/(2a) would become:

x = (-6 + 6) / 2 or x = (-6 -6) / 2

this would result in x = 0 or x = -6

one of the factors would be x.

the other factor would be x + 6

that agrees with what is shown.

i can't think of any other way to force the result to be in the form indicated.

i'll stick with my guess.

a = 10
r = -6
s = 0